OpenStudy (anonymous):

2x-y=6,3x+y=9.Solve it by evaluating proportional value of the variables.

OpenStudy (haseeb96):

2x-y=6...........eq#1 3x+y=9...........eq#2 take eq#1 y=2x-6......eq#3 put eq #3 into eq 2 3x +2x-6=9 5x =9+6 5x =15 x=15\5 x=3 now put the value of x into EQ#3 y=2(3)-6 y=6-6 y=0 So, th solution set is { ( 3, 0 ) }

OpenStudy (anonymous):

ur's is correct & it is solved by using substitution method ,but i need it by using elimination method......